# Julia in HEP

### v1

@@ Abstract
Julia is a mature general-purpose programming language, with a large ecosystem of libraries and more than 12000 third-party packages, which specifically targets scientific computing. As a language, Julia is as dynamic, interactive, and accessible as Python with NumPy, but achieves run-time performance on par with C/C++. In this paper, we describe the state of adoption of Julia in HEP, where momentum has been gathering over a number of years.
HEP-oriented Julia packages can already, via UnROOT.jl, read HEP’s major file formats, including TTree and RNTuple. Interfaces to some of HEP’s major software packages, such as through Geant4.jl, are available too. Jet reconstruction algorithms in Julia show excellent performance. A number of full HEP analyses have been performed in Julia.
@@ 1.1 HEP Needs
High-energy physics (HEP) is a large field, consisting of tens of thousands of researchers, almost all of whom will need to interact with software and contribute to software projects during their careers 2024EPJWC.29505023M . It is also one of the biggest, if not the biggest, generators of scientific datasets today, with exabytes of storage used by the LHC experiments Collaboration:2904204 . This data is processed by a huge corpus of software, estimated to be many tens of millions of lines in C++ hsfcwp .
This brings a challenge for HEP software. From the point of view of code efficiency we require fast execution, high throughput, and scalability at large computer centres and across distributed infrastructures. Considering human efficiency we would like a low barrier to entry for newcomers, the ability to prototype code rapidly, a broad ecosystem of well maintained packages, and excellent tooling for developers. These features are needed to make software able to deal efficiently with huge datasets, as well as accessible to a large group of developers.
@@ 2.1 Julia’s Motivations
Julia provides a syntax as productive as Python, especially for numerical work, whilst leveraging JAOT [^1] Just-Ahead-Of-Time compilation to provide speed similar to compiled, statically typed languages like C/C++ and Rust. It utilises type inference to allow coding in a “gradually typed” (§2.3.1), generic programming style, although the language will always track types for performance behind the scenes, and types can be specified explicitly if required. Like MatLab and Fortran, Julia’s native operations and type system support arrays as first-class entities, of any dimension, allowing array-oriented code to be productively generated and efficiently executed, with operations naturally “broadcast” element- or dimension-wise. This allows Julia to also be effective for writing, e.g., linear algebra heavy code. In addition, influence from the R community provides a wealth of statistical packages, and an R-flavoured approach to plots and visualisation. As with most languages of the 21st century, Julia is a fully open-source language, with its entire codebase freely available (and almost all of it written in Julia itself).
@@ 2.2 Julia in Practice
function mandel(z) c = z maxiter = 80 for n = 1:maxiter if abs2(z) > 4 return n-1 end z = z^2 + c end maxiter end
@@
Here we demonstrate importing of packages (line 1), function declaration with implicit types (line 3) and explicit loop (line 6) and branch constructs (line 7), implicit loops via “comprehensions” (line 15) to generate a value for every point in an implicitly defined array, and, finally, transparent broadcasting of operations over that array (line 16, where both the function call to Gray, and the division operation are distributed over the whole array). Unlike Python, explicit loops are optimised efficiently, and are no slower than comprehensions or “functional-style” iterators (which are also supported by Julia).
@@
The Julia Package “General Registry” indexes packages and their releases, and (like Rust and Javascript) relies on a public index hosted on GitHub. A help environment allows interactive help on any keyword or symbol known to the REPL; it is trivial to add documentation to a function or type by simply prepending its definition with a triple-quoted docstring.
@@
A well-supported VSCode extension is available for the language, which also supports the standard Language Server Protocol allowing it to support development in other editors. This includes the usual benefits of code completion (and Unicode completion for non-ASCII characters), linting, highlighting, and so on.
@@ 3.1 GPU Programming
Array based calculations are trivial to execute on the GPU. The listing in Figure 3 an array is copied to the GPU and then executes a trivial operation, but mainly shows that there is basically no boilerplate required to access GPU computations.
@@ 4.1 Challenges
C++ code is more difficult to interface to (a ‘feature’ of C++ itself), and the smoothest way to achieve it is to write a small wrapper in C++ to interface to Julia, the making use of the Julia interface package CxxWrap.jl. This process can be automated with the helper program WrapIt, which can generate wrapper code automatically from C++ headers. With this, wrapping very large libraries becomes much easier, as shown below in § 4.4.
@@ 4.2 HEP Data Formats
Most HEP experimental data is stored in ROOT ROOT:2011zz format, making it crucial for Julia to have the ability to read and write data in this format. Several options are available for achieving this functionality. One possibility is to create Julia bindings to the ROOT C++ package. However, this approach presents technical challenges, as it requires cross-compilation to generate binary artifacts, a process that is not yet fully supported by ROOT. Despite this, progress is being made with the ROOT.jl ROOT_jl package, which aims to address these challenges.
An alternative is the UnROOT.jl UnROOT:2022vnn package, which offers a pure Julia implementation of the ROOT I/O format, independent of ROOT or Python. This package supports transparent reading of both TTree and the newer RNTuple storage formats. While writing functionality is still under development, UnROOT.jl provides fast, memory-efficient, and lazy data access, reading only the necessary parts of a file on demand.
@@ 4.4 Simulation
Integrating Geant4 with Julia allows researchers to take advantage of Julia’s high-level programming capabilities and performance benefits while retaining the full functionality and efficiency of the Geant4 toolkit. This integration also provided an opportunity to rethink and streamline the interface, making it more intuitive and user-friendly. In particular, we focused on ensuring that application developers can concentrate on the essential aspects of their simulations while minimising configuration overhead. Boilerplate code and C++ idiosyncrasies are hidden, allowing for a cleaner, more concise approach to defining simulations. The performance of the Geant4.jl package is comparable to that of the native C++ Geant4 toolkit, demonstrating the feasibility of using Julia for HEP detector simulation.
@@ 4.5 Reconstruction
Reconstruction of HEP data is a core task for the field, and a huge amount of software, often rather detector specific, is dedicated to that task. Even for very detector specific code there are some early demonstrations that Julia would be suitable, and can be competitive with modern optimised C++ for a large experiment on multiple different GPU backends, e.g., for the CMS pixel detector. For smaller experiments, there are already end-to-end solutions based on Julia, as shown below in §4.7.
For generic reconstruction tasks there is less code in general, however, for the task of sequential jet reconstruction, the Fastjet package Cacciari:2011ma has seen very wide adoption across many experiments. There is now a native Julia reimplementation of many of the same algorithms as are found in Fastjet, in the JetReconstruction.jl package polyglot-jets-chep23 . Code ergonomics have been found to be superior, taking advantage of many of Julia’s features such as array broadcasting and generic programming. Performance is, on average 16% better than Fastjet for popular $pp$ reconstruction algorithms, and 33% better for the Durham algorithm for $e^{+}e^{-}$ events in FCCee studies fast-jet-reco-julia . Using Julia’s package extension mechanisms, support for the ubiquitous EDM4hep EDM4hep:2022leb event data model is supported directly. This adds to the user experience, where not only can the package offer simpler, more generic interfaces than Fastjet, but there is the ability to seamlessly hook into the rest of the Julia ecosystem, e.g., for jet visualisation with the native Julia Makie.jl package Danisch2021 .
@@ 5 Conclusions
As we have shown Julia is an ideal language for scientific computing. It combines runtime efficiency and scalability, which is as good as C++ and Fortran, with a modern high-productivity setup, suitable for rapid prototyping. The higher level features of the language lend themselves to compact code that expresses mathematical operations in a natural way. Multiple dispatch, supported by Julia’s type system, allows excellent code composition, with a very high degree of code reuse. Julia’s package system allows for easy setup of environments and excellent reproducibility. The language also interfaces with existing HEP code bases, allowing reuse of large codes, with improved ergonomics and integration with other Julia packages.

### v2

@@ Abstract
Julia is a mature general-purpose programming language, with a large ecosystem of libraries and more than 12000 third-party packages, specifically targeting scientific computing. Julia runs on x86, aarch64 and PowerPC architectures, and on all major GPU platforms. As a language, Julia is as dynamic, interactive, and accessible as Python with NumPy, but achieves run-time performance on par with C/C++. In this paper, we describe the state of adoption of Julia in HEP, where momentum has been gathering over a number of years.
HEP-oriented Julia packages can already read HEP’s major file formats, including TTree and RNTuple. Interfaces to some of HEP’s major software packages, such as Geant4, are available too. Jet reconstruction algorithms in Julia show excellent performance. A number of full HEP analyses have been performed in Julia.
@@ 1.1 HEP Needs
High-energy physics (HEP) is a large field, consisting of tens of thousands of researchers, almost all of whom will need to interact with software and contribute to software projects during their careers 2024EPJWC.29505023M . It is also one of the biggest, if not the biggest, generators of scientific datasets today, with exabytes of storage used by the LHC experiments Collaboration:2904204 . This data is processed by a huge corpus of software, estimated to be many tens of millions of lines in C++ and Python, as well as Fortran, CUDA and others hsfcwp .
This brings a challenge for HEP software. From the point of view of code efficiency we require fast execution, high throughput, and scalability at large heterogeneous computer centres and across distributed infrastructures. Considering human efficiency we would like a low barrier to entry for newcomers, the ability to prototype code rapidly, a broad ecosystem of well maintained packages, and excellent tooling for developers. These features are needed to make software able to deal efficiently with huge datasets, as well as accessible to a large group of developers. Stability and code preservation are also needed.
@@ 2.1 Julia’s Motivations
Julia provides a syntax as productive as Python, especially for numerical work, whilst leveraging JAOT [^1] Just-Ahead-Of-Time: meaning compilation to machine code happens, like C++, but just before a code block is executed, unlike C++.Julia can also cache these results between runs. compilation to provide speed similar to compiled, statically typed languages like C/C++ and Rust. It utilises type inference to allow coding in a “gradually typed” (§2.3.1), generic programming style, although the language will always track types for performance behind the scenes, and types can be specified explicitly if required. Like MatLab and Fortran, Julia’s native operations and type system support arrays as first-class entities, of any dimension, allowing array-oriented code to be productively generated and efficiently executed, with operations naturally “broadcast” element- or dimension-wise. This allows Julia to also be effective for writing, e.g., linear algebra heavy code. In addition, influence from the R community provides a wealth of statistical packages, and an R-flavoured approach to plots and visualisation. As with most languages of the 21st century, Julia is a fully open-source language, with its entire codebase freely available (and almost all of it written in Julia itself).
@@ 2.2 Julia in Practice
function mandel(z; maxiter=80) c = z for n = 1:maxiter if abs2(z) > 4 return n-1 end z = z^2 + c end maxiter end
@@
Here we demonstrate importing of packages (line 1); function declaration with implicit types, positional and named parameters (line 3); an explicit loop (line 5); branch constructs (line 6); the implicit Julia use of return values (line 11); implicit loops via “comprehensions” (line 14) to generate a value for every point in an implicitly defined array; and, finally, transparent broadcasting of operations over that array (line 15, where both the function call to Gray, and the division operation are distributed over the whole array). Unlike Python, explicit loops are optimised efficiently, and are no slower than comprehensions or “functional-style” iterators (which are also supported by Julia).
@@
The Julia Package “General Registry” indexes packages and their releases, and (like Rust and Javascript) relies on a public index hosted on GitHub. A help environment allows interactive help on any keyword or symbol known a the Read-Evaluate-Print-Loop (REPL) environment (which would be very familiar to Python users); it is trivial to add documentation to a function or type by simply prepending its definition with a triple-quoted docstring.
@@
A well-supported VSCode extension is available for the language, which also supports the standard Language Server Protocol allowing it to support development in other editors. This includes the usual benefits of code completion, including using Copilot and other AI extensions, Unicode completion, linting and highlighting. There is a built-in interactive debugger, as well as innovative tools to break into a REPL in the middle of code execution (Infiltrator.jl).
Julia is now popular enough to have many online training materials, an active support community on Slack and Discourse, and many postings on Stackoverflow.
@@ 3.1 GPU Programming
Array based calculations are trivial to execute on the GPU. In the listing in Figure 3 an array is copied to the GPU and then executes a trivial operation, but mainly shows that there is basically no boilerplate required to access GPU computations.
@@
Memory management on the GPU is mostly automated, including garbage collection of unreachable objects, but more fine grained steering is available.
@@ 4.1 Challenges
C++ code is more difficult to interface to (a ‘feature’ of C++ itself), and the smoothest way to achieve it is to write a small wrapper in C++ to interface to Julia, making use of the Julia interface package CxxWrap.jl. This process can be automated with the helper program WrapIt, which can generate wrapper code automatically from C++ headers. With this, wrapping very large libraries becomes much easier, as shown below in § 4.4.
@@ 4.2 HEP Data Formats
Most HEP experimental data is stored in ROOT ROOT:2011zz format, making it crucial for Julia to have the ability to read and write data in this format. Several options are available for achieving this functionality. One possibility is to create Julia bindings to the ROOT C++ package. However, this approach presents technical challenges, as it requires cross-compilation to generate binary artifacts, a process that is not yet fully supported by ROOT. Despite this, progress is being made with the ROOT.jl ROOT_jl package, which aims to address these challenges providing a direct interface to ROOT’s C++ code.
An alternative is the UnROOT.jl UnROOT:2022vnn package, which offers a pure Julia implementation of the ROOT I/O format, independent of ROOT or Python, thus implying a lightweight solution, with far fewer dependencies and a light. This package supports transparent reading of both TTree and the newer RNTuple storage formats. While writing functionality is still under development, UnROOT.jl provides fast, memory-efficient, and lazy data access, reading only the necessary parts of a file on demand.
@@ 4.4 Simulation
Integrating Geant4 with Julia allows researchers to take advantage of Julia’s high-level programming capabilities and performance (e.g., to plot results or code user actions), while retaining the full functionality and efficiency of the underlying Geant4 toolkit. This integration also provided an opportunity to rethink and streamline the interface, making it more intuitive and user-friendly. The package interface ensures that application developers can concentrate on the essential aspects of their simulations while minimising configuration overhead. Boilerplate code and C++ idiosyncrasies are hidden, allowing for a cleaner, more concise approach to defining simulations. The performance of the Geant4.jl package is comparable to that of the native C++ Geant4 toolkit, demonstrating the feasibility of using Julia for HEP detector simulation.
@@ 4.5 Reconstruction
Reconstruction of HEP data is a core task for the field, and a huge amount of software, often rather detector specific, is dedicated to that task. Even for very detector specific code there are some early demonstrations that Julia would be suitable, and can be competitive with modern optimised C++ for a large experiment on multiple different GPU backends, e.g., for the CMS pixel detector patatrack_julia . For smaller experiments, there are already end-to-end solutions based on Julia, as shown below in §4.7.
For generic reconstruction tasks there is less code in general, however, for the task of sequential jet reconstruction, the Fastjet package Cacciari:2011ma has seen very wide adoption across many experiments. There is now a native Julia reimplementation of many of the same algorithms as are found in Fastjet, in the JetReconstruction.jl package polyglot-jets-chep23 . Code ergonomics have been found to be superior, taking advantage of many of Julia’s features such as array broadcasting and generic programming. Performance is, on average 16% better than Fastjet for popular $pp$ reconstruction algorithms, and 33% better for the Durham algorithm for $e^{+}e^{-}$ events in FCCee studies fast-jet-reco-julia . Using Julia’s package extension mechanisms, the ubiquitous EDM4hep EDM4hep:2022leb event data model is supported directly. This adds to the user experience, where not only can the package offer simpler, more generic interfaces than Fastjet, but there is the ability to seamlessly hook into the rest of the Julia ecosystem, e.g., for jet visualisation with the native Julia Makie.jl package Danisch2021 .
@@ 5 Conclusions
As we have shown Julia is an ideal language for scientific computing. It combines runtime efficiency and scalability, which is as good as C++ and Fortran, with a modern high-productivity setup, suitable for rapid prototyping. AI and ML packages are supported, either as wrapped versions of Python/C++, or as native implementations. The higher level features of the language lend themselves to compact code that expresses mathematical operations in a natural way. Multiple dispatch, supported by Julia’s type system, allows excellent code composition, with a very high degree of code reuse. Julia’s package system allows for easy setup of environments and excellent reproducibility. The language also interfaces with existing HEP code bases, allowing reuse of large codes, with improved ergonomics and integration with other Julia packages.